### v1

@@ Abstract
In this short review, we provide an overview of recent progress in deploying advanced characterization techniques to understand the effects of local inhomogeneities in moiré heterostructures over multiple length scales. Particular emphasis is placed on correlating the impact of twist angle misalignment, nano-scale disorder, and atomic relaxation on the moiré potential and its collective excitations, particularly moiré excitons. Finally, we discuss future technological applications leveraging based on moiré excitons.
@@ I Introduction
In condensed matter, moiré patterns emerge when two or more atomically thin layers are artificially stacked with a slight angular misalignment or, in the case of heterolayers, are stacked featuring a small lattice constant mismatch Du et al. (2023b). Remarkably, the moiré interference patterns create a superstructure, the moiré lattice, with a new periodicity that is usually much larger than the underlying atomic lattices. This superlattice can dramatically alter the electronic band structure of the material, leading to the formation of mini-band structures and flat bands that directly affect the transport behaviour. Some key discoveries in this field of moiréengineering include the formation of flat bands Bistritzer and MacDonald (2011), promoting the relative relevance of correlations, but also the engineering of topological properties Choi et al. (2021), controlling polaritonsZhang et al. (2021) or networks of solitonsFu et al. (2020) and excitons, the key quasi-particle we will focus upon in our review. We refer the reader to recent reviews of moiré physics for an in-depth discussion of moiré physics:Behura et al. (2021); Alexeev et al. (2019); Mueller and Malic (2018); Huang et al. (2022); Ciarrocchi et al. (2022); Zhang et al. (2023a)
@@
Moiré excitons refer to photo-excited electron-hole bound states that form within (intra-) or between (inter-) layers with the potential defined by the moiré super-pattern Tran et al. (2019). Moiré excitons exhibit distinct energy levels and behaviors compared to their semiconductor-monolayer or three-dimensional counterparts. For a review of exciton physics in transition metal dichalcogenides and their applications, we refer the reader to Ref. Mueller and Malic (2018). Moiré excitons can host a rich tapestry of phenomena with valuable insights into novel quantum phenomena and many-body effects in low-dimensional systems, including exciton condensation, superfluidity, and topologically protected statesShimazaki et al. (2020); Choi et al. (2021). Moiré excitons also offer opportunities for tuning the electronic band structure of materials, enabling tailored properties for semiconductors, insulators, and topological materials, paving the way for technological innovationCiarrocchi et al. (2022). Additionally, they contribute a new dimension to valleytronics, allowing for control over valley polarization and dynamics in van der Waals heterostructuresUnuchek et al. (2019); Jin et al. (2019). Furthermore, moiré excitons play a pivotal role in realizing topological phases and the quantum anomalous Hall effect, leading to the development of innovative electronic and spintronic devicesLi et al. (2021b). Interactions between moiré excitons and antiferromagnetic ordering enable the manipulation of spin-dependent properties and spin information in optoelectronic devices Ramos et al. (2022). Moreover, moiré excitons present opportunities for controlling carrier dynamics and charge transport properties in heterostructures, resulting in high-performance optoelectronic devices with improved efficiency and reduced energy consumption. Excellent, complementary reviews of moiré excitons are Huang et al. (2022) focusing on early optical experiments and selection rules and Ciarrocchi et al. (2022) which highlights device architectures and engineering approaches to tailor excitons toward future applications.
@@
Controlling and understanding moiré superlattices, arising from the periodicity of moiré patterns, is fundamental to engineer moiré excitons and their dynamics phenomena Brem et al. (2020); Merkl et al. (2020); Meneghini et al. (2024). However, the excitonic properties of moiré heterostructures are extremely sensitive to local deviations from the ideal Lego-like description. In this short overview, we embark on a journey through multiple length scales and how new theoretical descriptions and advanced characterization techniques (Figure 1) have revealed the effects of twist angle misalignment, local inhomogeneities, and atomic relaxation in the physics of moiré excitons.
II Effect of spatial inhomogeneities on moiré heterostructures
@@
Excitonic states in moiré van der Waals superlattices offer exciting prospects for novel optoelectronics and quantum technologiesCiarrocchi et al. (2022). The strong exciton physics in TMD-based materials has been extensively investigated using different optical spectroscopy techniques, such as absorption spectroscopy, photoluminescence and photocurrent measurements, and optical pump-probe measurements Huang et al. (2022). While these techniques provide access to spectral and temporal degrees of freedom, they lack information about the real space properties of excitonic TMDs. However, local spatial modulations of the moiré superstructure due to strain fields, atomic defects Guo et al. (2021), and twist angle change extremely influence their optoelectronic properties. Correlating the averaged optical and transport properties with these local variations is needed to unambiguously resolve the role of the moiré energy and length scale. In the following, we review recent applications of advanced characterization techniques to disentangling emergent excitonic phenomena from moiré potential disorder in order of spatial resolution.
II.1 Determining the twist angle of moiré heterobilayers via non-linear optical techniques
@@
In the fabrication of heterobilayer devices, the twist angle, the effective misalignment between the main crystallographic directions of the two stack monolayers has emerged as a novel degree of freedom that controls the moiré potential and enables tunability of the optical emission and absorption spectrum of TMD heterobilayers Huang et al. (2022) (Fig. 2-(a)). For example, in closely aligned WS/MoSe heterobilayers, corresponding to (AA stacking) and (AB stacking), the almost perfect overlap between the two Brillouin zones leads to direct band gaps between inter- and intra-layer transitions. However, even small deviations from the AA or AB stacking lead to large changes in the band overlap in momentum space. Figure 2-(b) shows the evolution of the lower ( eV) and upper hybrid excitons ( eV) with twist angle as measured from reflection spectroscopyZhang et al. (2020a) revealing changes to the excitonic spectrum of more than meV as a function of the twist angle. Thus, it is fundamental to precisely determine the stacking angle between the heterobilayers for future devices based on moiré excitons.
In this context, the observation of strong non-linear responses dependent on the twist angle has emerged as a unique and versatile tool to characterize the formation of the moiré excitons Jin et al. (2019); Tran et al. (2019). Similarly, changes of symmetry and nonlinear responses associated by the interlayer coupling and charge transfer associated with the moiré potential have also been detected via non-linear responses Shree et al. (2021b); Deng et al. (2022); Yang et al. (2020). For strong electric fields, the interaction of light can be modeled by a resulting polarization given by:
@@
with as the linear and non-linear optical susceptibility tensors and as the permittivity of vacuum. These high-order processes are governed by the crystallographic symmetry and the electronic band structure, carrying fundamental information about the wavefunctions involved in the two-photon absorption processMa et al. (2021). Terms with an even number of powers on the electric field are only allowed in inversion symmetry-breaking compounds. The second-order term carries special importance as it is responsible of second-harmonic generation (SHG), in which the constructive interference of the two oscillating electric fields with frequency can give rise to a polarization density at twice the frequency Zhao et al. (2018). In a normal-incidence SHG experiment, a single wavelength or broadband light pulse with controlled polarization is focused onto the sample by a microscopic objective. Emission from the sample is first collimated by the objective and a short-pass filtered to eliminate contributions from the fundamental before being focused onto, a generally fixed, photodiode or CCD camera. The angular dependence of the intensity of the response with respect to the main crystallographic symmetries and the polarization of the incoming and outgoing beam enables the resolution of the underlying structural point group.
@@
For 2D transition metal dichalcogenide (TMD) monolayers, a strong SHG response with maxima along the armchair or zig-zag direction can be measured due to their inversion symmetry-breaking nature (point group ). When a heterobilayer is formed, the SHG signal from each layers coherently interferes resulting in a pattern that is extremely sensitive to small twist angle misalignment from perfect AA and AB stacking, as shown in Fig.2 Hsu et al. (2014); Psilodimitrakopoulos et al. (2021). Note that a twist angle of results in a completely destructive interference, fully suppressing the SHG response. Moreover, strong electric dipole SHG from centrosymmetric bilayer MoS with stacking has been observed Shree et al. (2021a); Zhang et al. (2023b) due to a resonance enhancement associated with inter- and intra-layer excitons, a phenomenon also expected to occur in heterobilayers hosting moiré excitons. Indeed, strong twist angle enhancement has been observed in MoS/WSe devices at the intralayer hybridized exciton energies Paradisanos et al. (2022). Comprehensive reviews of the enhancement of non-linear processes in twisted and heterostructures of TDMs can be found in Ref. Wen et al. (2019); Zhou et al. (2022).
@@
Optical contrast of the moiré heterobilayers beyond subwavelength resolution ( 20 nm) can be achieved by scanning near-field optical microscopy (SNOM) Vincent (2021). In a scattering SNOM (s-SNOM), the metallic AFM tip locally confines the electric light field. Thus, the excitonic spectrum of TMD heterostructures could be resolved by measuring the scattered light amplitude and phase as a function of position and energy McLeod et al. (2014). s-SNOM has been recently deployed to monolayers of MoSe2 and WSe2 and MoSe2/WSe2 heterobilayers to extract the excitonic resonance energy and interlayer hybridization with nanometer resolution Zhang et al. (2022). These early measurements have open the way for local measurements and control of moiré excitons. For example, the AFM tip from a SNOM has been used to introduce a vertical deformation in heterostructures of WSe/WS at cryogenic temperatures and concomitantly measure a shift on the moiré intralayer exciton as a function of uniaxial pressure Zhao et al. (2021), as shown in Fig. 2.
@@
Optical characterization techniques with real space imaging capabilities are thus fundamental to studying the properties of moiré heterostructure and excitons. Moreover, their large versatility enables a multi-messenger approach in which the combination of different wavelengths and scanning techniques characterize the excitation spectrum over a broad range of energies Sternbach et al. (2023).
@@
II.2 Nano-imaging of moiré exciton using photoemission microscopes
@@
Angle-resolved photoemission spectroscopy (ARPES) is a technique that explicitly accesses the material-specific electronic band structure Sobota et al. (2021). Its extension to the time domain (tr-ARPES) is sensitive to the transient dynamics of carriers with momentum resolution. In a tr-ARPES measurement, a pump pulse, typically in the visible or near-infrared range, excites electrons at a specific energy in the conduction band. Then, a second extreme ultraviolet probe pulse is applied to eject electrons from the material (Fig. 3 (a)) – according to Einstein’s photoelectric effect. Measuring the kinetic energy and the emission angle of the ejected electrons allows us to extract the energy and momentum of the initially excited electrons in the material Karni et al. (2023). Moreover, the successful development of high-repetition-rate extreme ultraviolet light sources pawed the way for tr-ARPES measurements on TMD materials. This powerful technique has been demonstrated to visualize momentum-dark excitons in TMD monolayers directly Madéo et al. (2020) and the momentum-resolved charge transfer dynamics in twisted TMD heterostructures Schmitt et al. (2022); Bange et al. (2024). While the first ARPES studies were performed on exciton dynamics in bulk TMD materials Bertoni et al. (2016); Wallauer et al. (2020); Dong et al. (2021); Grubišić Čabo et al. (2015), the more recent results, enabled by the development of time-of-flight momentum microscopes, have focused on the exciton physics of TMD monolayers Madéo et al. (2020); Wallauer et al. (2021); Kunin et al. (2023). Typically, tr-ARPES experiments do not spatially resolve the carrier dynamics, but they average over large sample areas of typically 10m or more. Thus, they cannot access spatial inhomogeneities, dielectric disorder, strain gradients, or local lattice reconstructions. Recently, ultrafast dark field momentum microscopy has been introduced as a new technique for simultaneous nano-imaging and nano-spectroscopy of materials Schmitt et al. (2023). The technique has a temporal resolution of 55 fs and a spatial resolution of nm and has been applied to study the spatial heterogeneity in a WSe-MoS heterostructure as well as its impact on the energy and the formation time of interlayer excitons. For a more in-depth discussion of the techniques’ details we refer the reader to other recent reviews Karni et al. (2023); Sobota et al. (2021).
@@
For excitonic materials, such as TMDs, tr-ARPES measurements provide access to exciton binding energies and their momentum in the reciprocal space. Here, the excitation pulse first creates excitons in the material. In contrast, the energy of the second pulse is used to break up the Coulomb-bound electron-hole pairs, photo-ejecting the electrons while leaving a hole density behind (Fig. 3 (a)). The ARPES signal of the photoemitted electrons reflects the excitonic band structure since the Coulomb-bound electrons and holes present a strongly correlated system Meneghini et al. (2023). The excitonic ARPES signal appears at the momentum corresponding to the ejected electron’s reciprocal-space valley. It is spectrally located one exciton binding energy below the conduction band. The momentum-resolved shape of the signal is characterized by the negative curvature of the valence band of the remaining hole. This negative dispersion from the photoemitted electron is the hallmark of the excitonic origin of the ARPES signal Christiansen et al. (2019); Man et al. (2021); Meneghini et al. (2023).
@@
In the following, we first review the exciton dynamics in TMD homo- and heterobilayers, which have attracted huge interest in current research. The latter exhibits a type-II band alignment favoring the tunneling of an electron or hole into the opposite layer and resulting in the formation of spatially separated interlayer excitons Yong et al. (2019), where the Coulomb-bound electrons and holes are located in different layers (Fig. 3 (b)-(c)). Finally, we discuss how the deployment of ultrafast dark field momentum microscopy has revealed local inhomogeneities of the dark excitons.
@@
ARPES signatures of the electron and hole bound in an interlayer exciton in a twisted WSe-MoS heterostructure have been observed [Fig. 3(d)-(i)] thanks to the high signal-to-noise ratio and high-quality samples with narrow initial valence band line widths Karni et al. (2022). The unexcited band structure of the heterostructure is characterized by two spin-split valence band maxima of the WSe layer at the K valley and a single broadband at lower energies corresponding to the valence band of the MoS layer (Fig. 3 (d)). After optically exciting the material resonant at the intralayer exciton in WSe, clear signatures of interlayer excitons about 1 eV above the valence band maximum are found (Fig. 4 (e)). Interestingly, the signal around the WSe valence band maximum becomes slightly depleted at the K and K’ points (Fig. 4 (f)). This was assigned to the holes created in WSe during the photoexcitation. Finally, the interlayer exciton signature at 1 eV exhibits a negative dispersion resembling the negative curvature of the WSe valence band (Fig. 3(i)). This anomalous dispersion has been theoretically predicted Christiansen et al. (2019) and is a hallmark of the excitonic origin of the observed ARPES signal.
Recently, it was shown that tr-ARPES is an ideal technique to visualize hybrid exciton states directly Meneghini et al. (2023). Tunneling between two TMD layers leads to a mixture of intra- and interlayer exciton states, resulting in hybrid excitons that inherit properties of both exciton species. Here, electrons and/or holes are in a superposition of both layers. The formation of such hybrid states is facilitated in naturally stacked TMD homobilayers, whereas the dominating excitons in TMD heterostructures are mostly of interlayer character. Thus, the performed study focuses on MoS homobilayers, where hybridized momentum-dark K excitons are the energetically lowest states. Here, the hole at the valley is strongly delocalized between the layers. During the ARPES process, the hybrid exciton breaks up into an ejected free electron from the K valley and a superposition of hybrid holes. This gives rise to a double-peak signal reflecting the superposition of the remaining holes between the two hybrid valence bands at the point.
@@
Figure 4 (a) illustrate how the momentum- and energy-resolved ARPES signal evolves in time. During the optical excitation, the spectrum is characterized by a pronounced single peak stemming from the optically excited intralayer exciton (Fig. 4 (a)). On a sub-100 fs timescale, these excitons scatter via emission of phonons to the energetically lower hybridized K excitons and the ARPES signal shows signatures of both KK intralayer and hybridized K excitons. After approx. 600 fs, an equilibrium distribution is reached with K excitons carrying nearly the entire population. Interestingly, their ARPES signature is characterized by two peaks, one slightly above the KK exciton and one red-shifted by more than 600 meV.

### v2

@@ Abstract
In this short review, we provide an overview of recent progress in deploying advanced characterization techniques to understand the effects of spatial variation and inhomogeneities in moiré heterostructures over multiple length scales. Particular emphasis is placed on correlating the impact of twist angle misalignment, nano-scale disorder, and atomic relaxation on the moiré potential and its collective excitations, particularly moiré excitons. Finally, we discuss future technological applications leveraging moiré excitons.
@@ I Introduction
The emergence of two-dimensional van der Waals (v.d.W.) materials has raised the intriguing idea of mixing and matching layers of different materials to realize novel properties in a given stack Geim and Grigorieva (2013). Moiré patterns emerge when two or more atomically thin v.d.W. layers are artificially stacked with a slight angular misalignment or, in the case of heterolayers, featuring a small lattice constant mismatch Du et al. (2023b). Remarkably, the moiré interference patterns create a superstructure, the moiré lattice, with a new periodicity usually much larger than the underlying atomic lattices. This superlattice can dramatically alter the electronic properties of the heterostructure. The long spatial periodicity of the moiré potential leads to the formation of minibands that directly modify the transport behavior. Some key discoveries in this field of moiré engineering include the formation of flat bands Bistritzer and MacDonald (2011), the enhancement of correlations Zhang et al. (2022b), the engineering of topological properties Choi et al. (2021), the formation and control of polaritons Zhang et al. (2021) or networks of solitons Fu et al. (2020). Moiré heterobilayers also enable the creation of excitons Jin et al. (2019b), the key quasi-particle we will focus upon in our review. We refer the reader to recent reviews for an in-depth discussion of moiré physics: Behura et al. (2021); Alexeev et al. (2019); Mueller and Malic (2018); Huang et al. (2022); Ciarrocchi et al. (2022); Zhang et al. (2023a); Liu et al. (2021). The mechanical properties of the v.d.W. compounds in the heterostructure are also profoundly impacted due to the moiré potential. While the effect of the moiré superlattice on the tribological properties of the interface is not the focus of this review, we refer the readers to the in-depth discussion presented in Yan et al. (2024).
@@
Moiré excitons refer to photo-excited electron-hole bound states that form within (intra-) or between (inter-) layers with the potential defined by the moiré superpattern Tran et al. (2019). Moiré excitons exhibit distinct energy levels and behaviors compared to their semiconductor-monolayer or three-dimensional counterparts. Moiré excitons can host a rich tapestry of phenomena with valuable insights into novel quantum phenomena and many-body effects in low-dimensional systems, including exciton condensation, superfluidity, and topologically protected states Shimazaki et al. (2020); Choi et al. (2021); Qi et al. (2023b). Moiré excitons also offer opportunities for tuning the electronic band structure of materials, enabling tailored properties for semiconductors, insulators, and topological materials, paving the way for technological innovation Ciarrocchi et al. (2022); Qi et al. (2022). Additionally, they contribute a new dimension to valleytronics, allowing for control over valley polarization and dynamics in v.d.W. heterostructures Unuchek et al. (2019); Jin et al. (2019a). Furthermore, moiré excitons play a pivotal role in realizing topological phases and the quantum anomalous Hall effect, leading to the development of innovative electronic and spintronic devices Li et al. (2021b). Interactions between moiré excitons and antiferromagnetic ordering enable the manipulation of spin-dependent properties and spin information in optoelectronic devices Ramos et al. (2022). Moreover, moiré excitons present opportunities for controlling carrier dynamics and charge transport properties in heterostructures, resulting in high-performance optoelectronic devices with improved efficiency and reduced energy consumption. Excellent, complementary reviews of moiré excitons are Huang et al. (2022); Regan et al. (2022) focusing on early optical experiments and selection rules. Moreover, Ciarrocchi et al. (2022) and Mueller and Malic (2018) highlight device architectures and engineering approaches to tailor excitons toward future applications.
@@
Controlling and understanding moiré superlattices, arising from the periodicity of moiré patterns, is fundamental to engineer moiré excitons and their dynamics Brem et al. (2020); Merkl et al. (2020); Meneghini et al. (2024). However, the excitonic properties of moiré heterostructures are extremely sensitive to local deviations from perfect twist angles and Lego-like stacking description. In this short overview, we discuss how new theoretical descriptions and advanced characterization techniques have been deployed to study the role of spatial variations over multiple length scales, namely twist angle misalignment, local inhomogeneities, and atomic relaxation, affect the moiré potential and the physics of moiré excitons. We conclude by discussing possible technological applications based on integrable moiré devices that motivate further efforts to understand and control local spatial variations.
II Characterizing the effects of spatial variation on the moiré potential and moiré excitons by advanced tools.
@@
Excitonic states in moiré v.d.W. superlattices offer exciting prospects for novel optoelectronics and quantum technologies Ciarrocchi et al. (2022). The strong exciton physics in transition metal dichalcogenides (TMDs) has been extensively investigated using different optical spectroscopy techniques, such as absorption spectroscopy, photoluminescence and photocurrent measurements, and optical pump-probe measurements Huang et al. (2022). While these techniques provide access to spectral and temporal degrees of freedom, they lack information about the real space properties of excitonic TMDs. However, local spatial modulations of the moiré superstructure due to strain fields, atomic defects Guo et al. (2021), and twist angle change extremely influence their optoelectronic properties. Correlating the averaged optical and transport properties with these local variations is needed to unambiguously resolve the role of the moiré energy and length scale.
In the following, we review recent applications of advanced characterization techniques to disentangling emergent excitonic phenomena from moiré potential disorder in order of spatial resolution. In particular, as highlighted in Figure 1, we focus on Second Harmonic Generation (SHG), momentum microscopy, atomic force microscopy (AFM) based techniques, and laser Scanning Tunneling Microscopy (STM). In Section II.1, we discuss how the extreme sensitivity of SHG, a non-linear optical technique, to inversion symmetry breaking enables quick and non-intrusive detection of deviations from the twist angle in moiré heterostructures Psilodimitrakopoulos et al. (2021). We also review how scattering Scanning Near-field Optical Microscopy (s-SNOM), a technique that combines the spatial resolution of AFM with optical spectroscopy, has been deployed to study variations in the excitonic response with nanometer resolution Zhang et al. (2022a). In Section II.2, we review the role of time-resolved Angle Resolved Photoemission Spectroscopy (tr-ARPES) in characterizing the formation dynamics of hybrid moiré excitons Meneghini et al. (2023). We also discuss how ultrafast dark field momentum microscopy, a recently introduced technique, reveals the role of local spatial inhomogeneities in the charge transfer dynamics of moiré excitons Schmitt et al. (2023, 2022). Finally, in Subsection II.3, we discuss how piezoresponse force microscopy (PFM), an extension of AFM, and STM have been deployed to study the moiré potential and deviations from the ideal membrane-like behavior of the layers Halbertal et al. (2022). Moreover, the extension of STM to the time domain by its combination with laser pulses Mogi et al. (2022) has revealed the role of corrugation and strain in the lifetime of moiré excitons.
II.1 Non-linear optical techniques deployed to moiré heterostructures.
@@
In the fabrication of heterobilayer devices, the twist angle, , the effective misalignment between the main crystallographic directions of the two stacked monolayers, has emerged as a novel degree of freedom that controls the moiré potential and enables tunability of the optical emission and absorption spectrum of TMD heterobilayers Huang et al. (2022). In general, perfect stacking can be achieved at (AA stacking) and (AB stacking). The almost perfect overlap between the Brillouin zones of each monolayer can lead moiré inter- and intra-layer excitons. The onset of excitonic correlations depends on electrostatic doping, on the v.d.W. materials participating in the heterostructure, and on twist angle Tang et al. (2020); Barré et al. (2022). For example, perfectly stacked heterostructures of WSe2/WS2 display more strongly interacting hybrid excitons than those of MoSe2/WS2 due to differences in the local TMD dielectric constants Tang et al. (2020). Small deviations from the AA or AB stacking lead to large changes in the band overlap in momentum space and a reconfiguration of the excitonic energies. This is illustrated in Figure 2 (a), where reflection spectroscopy measurements of the lower ( eV) and upper hybrid excitons ( eV) revealed changes to the excitonic spectrum of more than meV as a function of the twist angle Zhang et al. (2020a). Optical absorption spectroscopy has shown a similar dependence of the excitonic spectrum on the twist angle in heterobilayers of WSe2/MoSe2 Barré et al. (2022). Thus, precisely determining the stacking angle between the heterobilayers is fundamental for future devices based on moiré excitons. In this context, the observation of strong non-linear responses dependent on the twist angle has emerged as a unique and versatile tool to characterize the formation of the moiré excitons Jin et al. (2019a); Tran et al. (2019). Similarly, changes of symmetry and nonlinear responses associated with the interlayer coupling and charge transfer as a result of the moiré potential have also been detected via non-linear responses Shree et al. (2021b); Deng et al. (2022); Yang et al. (2020).
For strong electric fields, the interaction of light can be modeled by a resulting polarization given by:
@@
with as the linear and non-linear optical susceptibility tensors and as the permittivity of vacuum. These high-order processes are governed by the crystallographic symmetry and the electronic band structure, carrying fundamental information about the wavefunctions involved in the two-photon absorption process Ma et al. (2021). Terms with an even number of powers on the electric field are only allowed in inversion symmetry-breaking compounds. The second-order term carries special importance as it is responsible for second-harmonic generation (SHG), in which the constructive interference of the two oscillating electric fields with frequency can give rise to a polarization density at twice the frequency Zhao et al. (2018). In a normal-incidence SHG experiment, a microscopic objective focuses a single wavelength or broadband light pulse with controlled polarization onto the sample. Emission from the sample is first collimated by the objective and short-pass filtered to eliminate contributions from the fundamental beam before being focused onto a photodiode or CCD camera. Resolving the underlying structural point group can then be achieved by tracking the angular dependence of the intensity of the response with respect to the main crystallographic symmetries and the polarization of the incoming and outgoing beam.
@@
For 2D TMD monolayers, a strong SHG response with maxima along the armchair or zig-zag direction can be measured due to their inversion symmetry-breaking nature (point group ). When a heterobilayer is formed, the SHG signal from each layer coherently interferes, resulting in a pattern that is extremely sensitive to small twist angle misalignment from perfect AA and AB stacking, as shown in Figure 2 (b) Hsu et al. (2014); Psilodimitrakopoulos et al. (2021). Note that a twist angle of results in completely destructive interference, fully suppressing the SHG response, while it is maximal for . Moreover, strong electric dipole SHG from centrosymmetric bilayer MoS2 with stacking has been observed Shree et al. (2021a); Zhang et al. (2023b) due to a resonance enhancement associated with inter- and intra-layer excitons, a phenomenon also expected to occur in heterobilayers hosting moiré excitons. Indeed, strong twist angle enhancement has been observed in MoS2/WSe2 devices at the intralayer hybridized exciton energies Paradisanos et al. (2022). This is highlighted in Figure 2 (c). The SHG response is dominated by two main resonances at twice the energy of the A1s exciton of the MoSe2 (1.92 eV) and WSe2 (1.72 ev) monolayers. At those energies, the SHG polar plot at the heterostructure selectively reflects that of the corresponding monolayer. Comprehensive reviews of the enhancement of non-linear processes in twisted and heterostructures of TMDs can be found in Ref. Wen et al. (2019); Zhou et al. (2022). Thus, SHG has emerged as a key technique in the study of moiré heterostructures to confirm the desired twist angle quickly, versatilely, and non-disruptively. However, the extreme surface sensitivity of SHG requires clean and flat samples. Moreover, SHG is limited by the diffraction limit of light, making its deployment to small and complex stacks challenging.
@@
Optical contrast of the moiré heterobilayers beyond subwavelength resolution ( 20 nm) can be achieved by scanning near-field optical microscopy (SNOM) Vincent (2021). In a scanning SNOM (s-SNOM), the metallic AFM tip locally confines the electric light field. Thus, the excitonic spectrum of TMD heterostructures could be resolved by measuring the scattered light amplitude and phase as a function of position and energy McLeod et al. (2014). s-SNOM has been recently deployed to monolayers of MoSe2 and WSe2 and MoSe2/WSe2 heterobilayers to extract the excitonic resonance energy and interlayer hybridization with nanometer resolution Zhang et al. (2022a) [Figure 2 (d)-(e)]. However, s-SNOM is not free of sample-tip interactions that can modify the properties of the heterostructure, which can be deployed to control the electronic structure of moiré excitons. For example, the AFM tip from a SNOM has been used to introduce a vertical deformation in heterostructures of WSe2/WS2 at cryogenic temperatures and concomitantly measure a shift on the moiré intralayer exciton as a function of uniaxial pressure Zhao et al. (2021), as shown in Figure 2 (f). The exciton energy reduces by more than meV for vertical displacements that correspond to GPa.
@@
Optical characterization techniques with real space imaging capabilities are thus fundamental to studying the properties of moiré heterostructure and excitons. Moreover, their large versatility enables a multi-messenger approach combining different wavelengths and scanning techniques to characterize the excitation spectrum over a broad range of energies Sternbach et al. (2023).
@@
II.2 Real and reciprocal space imaging of moiré excitons using photoemission microscopes
@@
Angle-resolved photoemission spectroscopy (ARPES) can directly access a material-specific electronic band structure Sobota et al. (2021). In an ARPES measurement, an extreme ultraviolet probe pulse is applied to eject electrons from the material [Figure 3 (a)]. Measuring the kinetic energy and the emission angle of the ejected electrons allows us to extract the energy and momentum of the initially excited electrons in the material Karni et al. (2023). The successful development of high-repetition-rate extreme ultraviolet light sources pawed the way for time-resolved ARPES measurements (tr-ARPES) on TMD materials. In a tr-ARPES experiment, a second pump pulse, typically in the visible or near-infrared range, enables sensitivity to the transient dynamics of photoexcited carriers with momentum resolution. For excitonic materials, such as TMDs, tr-ARPES measurements provide access to exciton binding energies and their momentum in the reciprocal space. This powerful technique has been demonstrated to directly visualize momentum-dark excitons in TMD monolayers Madéo et al. (2020) and the momentum-resolved charge transfer dynamics in twisted TMD heterostructures Schmitt et al. (2022); Bange et al. (2024). While the first ARPES studies were performed on exciton dynamics in bulk TMD materials Bertoni et al. (2016); Wallauer et al. (2020); Dong et al. (2021); Grubišić Čabo et al. (2015), more recent results have focused on the exciton physics of TMD monolayers enabled by the development of time-of-flight momentum microscopes Madéo et al. (2020); Wallauer et al. (2021); Kunin et al. (2023). For a more in-depth discussion of the techniques’ details, we refer the reader to other recent reviews Karni et al. (2023); Sobota et al. (2021).
@@
We start by discussing how tr-ARPES can reveal the dynamics of excitons in TMDs with momentum resolution. The pump (excitation) pulse is tuned to create excitons in the material. The second pulse breaks up the Coulomb-bound electron-hole pairs, photo-ejecting the electrons while leaving a hole density behind [Figure 3 (a)]. The ARPES signal of the photoemitted electrons reflects the excitonic band structure since the Coulomb-bound electrons and holes act as a strongly correlated system Meneghini et al. (2023). In energy, the ARPES response is spectrally located one exciton binding energy below the conduction band. The excitonic ARPES signal appears at the momentum point corresponding to the ejected electron’s reciprocal-space valley, and its dispersion is characterized by the negative curvature of the valence band of the remaining hole Christiansen et al. (2019); Man et al. (2021); Meneghini et al. (2023).
@@
The rest of this section is organized as follows. We first review the exciton dynamics in TMD homo- and heterobilayers. In the latter case, type-II band alignment favors the tunneling of an electron or hole into the opposite layer. This leads to spatially separated interlayer excitons Yong et al. (2019), where the Coulomb-bound electrons and holes are located in different layers [Figure 3 (b)-(c)]. Second, we introduce ultrafast dark field momentum microscopy a new technique for simultaneous nano-imaging and nano-spectroscopy of materials Schmitt et al. (2023) with a temporal resolution of 55 fs and a spatial resolution of nm. Finally, we discuss how the deployment of ultrafast dark field momentum microscopy has revealed local inhomogeneities of the dark excitons.
@@
In a twisted WSe2-MoS2 heterostructure, ARPES signatures of the bound electron and hole pairs in an interlayer exciton have been observed [Figure 3 (d)-(i)]. This was enabled by the high signal-to-noise ratio and high-quality samples with narrow initial valence band line widths Karni et al. (2022). The unexcited band structure of the twisted WSe2-MoS2 heterostructure is characterized by two spin-split valence band maxima of the WSe2 layer at the K valley and a single broad valence band of the MoS2 layer [Figure 3 (d) and (f)]. After optically exciting the material with photons resonant at the WSe2 intralayer exciton, clear signatures of interlayer excitons about 1 eV above the valence band maximum are found [Figure 3 (e)]. Concomitantly, the signal around the WSe2 valence band maximum becomes slightly depleted at the K and K’ points [Figure 3 (g) and (h)]. This was assigned to the creation of holes in WSe2 during photoexcitation. Finally, the interlayer exciton signature at 1 eV exhibits a negative dispersion resembling the negative curvature of the WSe2 valence band [Figure 3(i)]. This anomalous dispersion has been theoretically predicted Christiansen et al. (2019) and is a hallmark of the excitonic origin of the observed ARPES signal.
Recently, it was shown that tr-ARPES is an ideal technique to directly visualize hybrid exciton states in moiré heterostructures Meneghini et al. (2023). Hybrid moiré excitons refer to excitonic states formed by a mixture of intra- and interlayer exciton states due to the direct tunneling between two TMD layers. In other words, electrons and holes are in superposition in both layers of the moiré heterostructures. The formation of such hybrid excitonic states is facilitated in naturally stacked TMD homobilayers, making them the dominating excitons of interlayer character as first observed in WSe2/WS2 moiré heterostructures Naik et al. (2022).
@@
Thus, we first review a tr-ARPES study in MoS2 homobilayers Meneghini et al. (2023). Figure 4 (a) illustrates how the momentum- and energy-resolved ARPES signal evolves in time. During the optical excitation, the spectrum is predicted to be characterized by a pronounced single peak stemming from the optically excited intralayer exciton [Left panel of Figure 4 (a)]. On a sub-100 fs timescale, these excitons scatter via emission of phonons to the energetically lower hybridized -K excitons and the ARPES signal shows signatures of both K-K intralayer and hybridized -K excitons [Middle panel of Figure 4 (a)]. After fs, an equilibrium distribution is reached with -K excitons carrying nearly the entire population [Right panel of Figure 4 (a)]. The ARPES signal is predicted to display two peaks, one slightly above the K-K exciton and one red-shifted by more than 600 meV. This work revealed that the energetically lowest states are the hybridized momentum-dark -K excitons Meneghini et al. (2023). Here, the hole at the valley is strongly delocalized between the layers. During the ARPES process, the hybrid exciton breaks up into an ejected free electron from the K valley and a superposition of hybrid holes. This gives rise to a double-peak signal reflecting the superposition of the remaining holes between the two hybrid valence bands at the point.